# Gromov’s Compactness Theorem for the Intrinsic Timed Hausdorff Distance

### v1

@@ Abstract.
We prove Gromov’s Compactness Theorem for the Intrinsic timed Hausdorff ($\tau-H$) convergence of timed-metric-spaces using timed-Fréchet maps. Our proof introduces the notion of ”addresses” and provides a new way of stating Gromov’s original compactness theorem for Gromov-Hausdorff (GH) convergence of metric spaces. We also obtain a new Arzela-Ascoli Theorem for real valued uniformly bounded Lipschitz functions on GH converging compact metric spaces. The intrinsic timed-Hausdorff distance between timed-metric-spaces was first defined by Sakovich-Sormani to define a weak notion of convergence for space-times, and our compactness theorem will soon be applied to advance their work in this direction.
@@ 1. Introduction
One of the most profound theorems in metric geometry is Gromov’s Compactness Theorem: a sequence of equibounded and equicompact metric spaces has a subsequence converging in the Gromov-Hausdorff sense to a compact metric space. See Gromov’s presentation in [16] and his application of a stronger version to groups of polynomial growth in [15]. We review these notions and state this theorem precisely in Sections 2.12.3.
@@
When a sequence of metric spaces converges in the Gromov-Hausdorff sense then one can control distances on the limit spaces. Not only can one control distances on these converging sequences of metric spaces, but one can also prove Arzela-Ascoli theorems for sequences of uniformly bounded Lipchitz functions on the sequence of spaces (c.f. Greene-Petersen [14]).
@@
Before Gromov’s Compactness Theorem, Riemannian Geometers could only study the limits of sequences of diffeomorphic Riemannian manifolds. With Gromov-Hausdorff convergence, geometers can study sequences of Riemannian manifolds with different topologies and even different dimensions. A sequence of compact Riemannian manifolds can be canonically converted into compact metric spaces using the Riemannian distance. With uniform upper diameter bounds and lower Ricci curvature bounds, Gromov proved that Riemannian manifolds are equicompact and equibounded and thus any sequence subconverges to a limit space [16][17]. This was applied with great success by Cheeger, Colding, Fukaya, Naber, Sormani, Wei and others in many papers including [10][13][8][9][30][11].
@@
Lorentzian geometers would also like to study sequences of smooth Lorentzian manifolds which are not diffeomorphic. In [25], Sakovich and Sormani introduced the notion of the intrinsic timed-Hausdorff convergence as one possible way to achieve this goal. They canonically converted a sequence of smooth causally-null-compactifiable space-times into timed-metric-spaces, $(X_{j},d_{j},\tau_{j})$, which are metric spaces, $(X_{j},d_{j})$, endowed with bounded Lipschitz one time functions, $\tau_{j}:X_{j}\to[0,\tau_{max}]$. They then defined the intrinsic timed-Hausdorff convergence of these timed-metric-spaces. We review these definitions in Sections 2.52.7.
In this paper we prove a compactness theorem for this notion of timed-Hausdorff convergence: a sequence of equibounded and equicompact timed-metric-spaces with uniformly bounded time functions converges in the intrinsic timed-Hausdorff sense. For the full somewhat stronger statement, see Theorem 3.1.
Our compactness theorem is proven in Sections 3.13.3. In Section 3.1 we begin the proof by imitating Gromov’s proof of his compactness theorem in [15]. As we added significant details to his original proof, we developed the notion of “addresses” for points in Section 3.2. With these “addresses” we complete a stronger statement of the compactness theorem which controls the convergence of time in a uniform way in Section 3.3.
Our notion of “addresses” provides a new way of stating Gromov’s Compactness Theorem for sequences of metric spaces, that we state and prove in Section 4.1. We also restate an Arzela-Ascoli Theorem for sequences of real valued uniformly bounded Lipschitz functions on a converging sequence of metric spaces using “addresses” in Section 4.2. We suggest a more general Arzela-Ascoli Conjecture that could be proven using “addresses” in Section 4.3.
Sakovich-Sormani introduced the notion of the intrinsic timed-Hausdorff distance in [25] to define a new notion of convergence for smooth Lorentzian space-times. They describe a method of converting smooth causally null compactifiable space-times into compact timed metric spaces in a canonical way using the cosmological time function of Anderson-Galloway-Howard [4] and the null distance of Sormani-Vega [29]. Although we do not wish to go into the details of this conversion in this paper, we point out that our compactness theorem is immediately a precompactness theorem for sequences of such smooth Lorentzian manifolds producing limit spaces that are timed-metric-spaces in [25]. Further applications of our compactness theorem to Lorentzian manifolds will appear in future work.
It should be noted that some people have used the Sormani-Vega null distance of [29] to convert space-times into metric spaces, without keeping track of the time functions. See the work of Allen, Burtscher, García-Heveling, Kunzinger, and Sämann in [2, 3, 6, 19]. If one does this, then one can apply Gromov’s original compactness theorem to obtain a timeless GH limit space which is only a metric space. However, one needs a time function to define the causal structure on the limit space [25]. We believe the work of Allen, Burtscher, and García-Heveling should be extendable to the stronger notion of timed-Hausdorff convergence studied here.
@@
There are special classes of space-times where the cosmological time functions are easy to control. These include the big-bang space-times and future developed space-times. Sormani-Vega [28] and Sakovich-Sormani [25] defined convergence for such space-times based on pointed GH convergence of Gromov [17] and GH Convergence for Metric Pairs introduced by Che, Galaz-García, Guijarro, and Membrillo Solis in [7]. Such notions already have compactness theorems proven by Gromov [17] and by Che and Ahumada Gómez [1] respectively. Sakovich-Sormani defined the intrinsic timed-Hausdorff distance to apply to a far wider class of space-times including sub-space-times that exhaust asymptotically flat space-times [25] and we need our new compactness theorem to apply in these settings.
@@
Other notions of weak convergence of space-times do not involve the convergence of the space-times into metric spaces or timed-metric-spaces. Minguzzi-Suhr [21] have a compactness result in the setting of “Lorentzian metric spaces” (which are not metric spaces) building upon work of Noldus [23, 24]. See also the work of Mondino–Sämann [22] in the setting of “Lorentzian pre-length spaces”, as introduced in [18], where they define a notion of Lorentzian Gromov–Hausdorff convergence and obtain a precompactness principle using causal diamonds and Lorentz distances. The spaces considered in these works are not metric spaces. It is not yet clear how their notions of convergence relate to the Sakovich-Sormani notions defined using timed-metric-spaces in [25]. This is worth exploring further.
@@
Finally we note that, in upcoming work, Sakovich-Sormani will be introducing timed intrinsic flat convergence of space-times whose limit spaces are rectifiable timed-metric-spaces with biLipchitz charts [26]. Our compactness theorem proven in this paper will be applied in that paper in an essential way.
@@
Acknowledgements: We would like to thank Anna Sakovich (Uppsala) for online conversations in July 2025 that lead us to pursuing this project. We all gratefully acknowledge support from the Simons Center for Geometry and Physics, Stony Brook University where we had the opportunity to meet together in person in September 2025 as part of the Program Geometry and Convergence in Mathematical General Relativity. R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. C. Sormani was funded in part by a PSC-CUNY grant. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ 2.1. Review of the Hausdorff Distance
Recall that the Hausdorff distance between sets, $U,W\subset Z$, is extrinsic in that it depends on the distance in the extrinsic space:
The Hausdorff distance between subsets, $U,W\subset Z$, lying in a common metric space, $(Z,d_{Z})$, id defined
(1) $d_{H}^{Z}\bigg(U,W\bigg)=\inf\left\{r>0:\begin{array}[]{c}\forall u\in U,\exists w\in W\textrm{ s.t. }d_{Z}(u,w)<r\\ \forall w\in W,\exists u\in U\textrm{ s.t. }d_{Z}(u,w)<r\end{array}\right\}.$
2.2. Review of Gromov-Hausdorff Distance
Gromov defined his intrinsic Hausdorff (GH) distance between distinct metric spaces which do not lie in a common metric space in [17]. This is now called the Gromov-Hausdorff distance. The GH distance is intrinsic in the sense that it depends only on the intrinsic geometry of the two given metric spaces, $(X_{j},d_{j})$, and not the extrinsic geometry of some common ambient space.
In order to define this distance intrinsically, Gromov considered all possible common metric spaces, $(Z,d_{Z})$, and applied maps: $\varphi_{j}:X_{j}\to Z$ that are distance preserving
@@ 2.2. Review of Gromov-Hausdorff Distance
Thanks to the infimum taken over all possible distance preserving maps, the Gromov-Hausdorff distance depends only on the intrinsic geometry of the original pair of metric spaces not on any particular external metric space, $(Z,d_{Z})$. Yet at the same time the GH distance between metric spaces has many of the same properties as the Hausdorff distance. In particular there are correspondences between $X_{1}$ and $X_{2}$ with controlled distance distortion that are often used to define the GH distance, but we will not use correspondences in this paper.
@@ 2.3. Review of Gromov’s Compactness Theorem
Recall Gromov’s Compactness Theorem: proven in [17] and [15]:
@@
If $(X_{j},d_{j},\tau_{j})$ are a sequence of compact metric spaces that are equibounded,
@@
then a subsequence converges in the Gromov-Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty})$:
@@
In fact, there exist a uniform compact metric space, $(Z,d_{Z})$, and distance preserving maps,
(7) $\varphi_{j}:X_{j}\to Z,$
@@ 2.4. Review of Fréchet Maps
Here we review the notion of a Fréchet map [12].  [^1] In [25] this was accidentally attributed to Kuratowski [20] and will be fixed.
First recall that a separable metric-space, $(X,d)$, is a metric-space containing a countable dense collection of points, $\{x_{1},x_{2},...\}\subset X$.
@@
Given a separable metric-space, $(X,d)$, for any countable dense collection of points $\mathcal{N}=\{x_{1},x_{2},$...\} of $X$, one can define a Fréchet map
(11) $\kappa_{X}=\kappa_{X,\mathcal{N}}:(X,d)\to(\ell^{\infty},d_{\ell^{\infty}})$
@@
[…]
 [^2] In v1 of [25] this was accidentally attributed to Kuratowski [20] and will be fixed.
The following theorem has been credited to Fréchet:
@@
It is well known that one can define a version of the Gromov-Hausdorff distance by taking the infimum of the Hausdorff distance between images of Fréchet maps into $\ell^{\infty}$.
@@
The Fréchet-GH distance
@@
which is an infimum over all selections of countably dense points and reorderings of these selected points in $X$ and in $Y$.
By Fréchet’s Theorem, one immediately has
@@
In fact this notion is biLipschitz equivalent to the Gromov-Hausdorff distance:
(17) $d_{GH}((X,d_{X}),(Y,d_{Y}))\leq d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))\leq 2d_{GH}((X,d_{X}),(Y,d_{Y}))$
for any pair of compact metric-spaces. This close relationship with the Gromov-Hausdorff distance is discussed and reproven in the background of [25] and directly inspired their definition of the intrinsic timed Hausdorff distance using timed-Fréchet maps.
@@ 2.5. Review of Timed-Metric-Spaces
The notion of a timed-metric-space was introduced by Sakovich-Sormani in [25] because they had constructed a canonical way to convert certain classes of smooth Lorentzian space-times into timed-metric-spaces keeping track of the causal structure of the original Lorentzian manifold. For our purposes, we need only the abstract notion of a timed-metric-space:
@@
A timed-metric-space, $(X,d,\tau)$, is a metric space, $(X,d)$, with a definite, symmetric map $d:X\times X\to[0,\infty),$ satisfying the triangle inequality, that is endowed with a Lipschitz one time function, $\tau:X\to[0,\infty).$ It has a causal structure defined as follows:
(18) $p\in J^{+}(q)\Longleftrightarrow\tau(p)-\tau(q)=d(p,q).$
Note that a map $F:(X,d_{X},\tau_{X})\to(Y,d_{Y},\tau_{Y})$ is distance preserving iff
@@
It preserves the causal structure iff
@@
In this paper we focus on the geometry of timed-metric-spaces and the intrinsic timed Hausdorff distance. Our work applies to study Lorentzian manifolds and their intrinsic timed Hausdorff limits, but we will not be exploring that relationship within this paper. We will not refer further to smooth manifolds.

### v2

# Gromov’s Compactness Theorem for the Intrinsic Timed-Hausdorff Distance

@@ Abstract.
The intrinsic timed-Hausdorff distance between timed-metric-spaces, first introduced by Sakovich–Sormani, yields a weak notion of convergence for space-times. In this paper we prove a compactness theorem for the intrinsic timed-Hausdorff convergence of timed-metric-spaces using timed-Fréchet maps. Our proof introduces the notion of “addresses” and provides a new way of stating Gromov’s original compactness theorem for Gromov–Hausdorff convergence of metric spaces. We also obtain a new Arzelà–Ascoli theorem for real valued uniformly bounded Lipschitz functions on Gromov–Hausdorff converging compact metric spaces.
@@ 1. Introduction
One of the most profound theorems in metric geometry is Gromov’s Compactness Theorem, which states that a sequence of equibounded and equicompact metric spaces has a subsequence converging in the Gromov–Hausdorff sense to a compact metric space [16, 17]. Moreover, when a sequence of metric spaces converges in the Gromov–Hausdorff sense one can prove Arzelà–Ascoli theorems for sequences of uniformly bounded Lipchitz functions on the sequence of spaces (c.f. Greene–Petersen [15]).
@@
A sequence of compact Riemannian manifolds can be canonically converted into compact metric spaces using the Riemannian distance. With Gromov–Hausdorff convergence, one can study sequences of Riemannian manifolds with different topologies and even different dimensions. With uniform upper diameter bounds and lower Ricci curvature bounds, Gromov proved that Riemannian manifolds are equicompact and equibounded and thus any sequence subconverges to a limit space [17, 18]. This was applied with great success by Cheeger, Colding, Fukaya, Naber, Sormani, Wei and others in many papers including [8, 9, 10, 11, 14, 31].
@@
One would also like to study sequences of smooth Lorentzian manifolds which are not diffeomorphic. In [26], Sakovich and Sormani introduced the notion of the intrinsic timed-Hausdorff convergence as one possible way to achieve this. They described a method of converting smooth causally null compactifiable space-times into compact timed-metric-spaces in a canonical way using the cosmological time function of Anderson–Galloway–Howard [4] and the null distance of Sormani–Vega [30]. In general, timed-metric-spaces, $(X_{j},d_{j},\tau_{j})$, are metric spaces, $(X_{j},d_{j})$, endowed with a Lipschitz one  [^1] In this manuscript we refer to Lipschitz functions with Lipschitz constant bounded above by $1$ as Lipschitz one function, $\tau_{j}\colon X_{j}\to\mathbb{R}$, and a causal structure. They then defined the intrinsic timed-Hausdorff convergence of these timed-metric-spaces.
In this paper we prove a compactness theorem for the notion of intrinsic timed-Hausdorff convergence. For the full, somewhat stronger statement, see Theorem 3.1. Note that no knowledge of Lorentzian geometry is required throughout this manuscript, but further applications of our compactness theorem to Lorentzian manifolds will appear in future work.
Theorem 1.1 (Timed Gromov Compactness Theorem–short version).
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed metric spaces that are equibounded,
(1) $\exists D>0\,s.t.\,\operatorname{diam}_{d_{j}}(X_{j})\leq D,$
equicompact,
and have uniform bounds on their Lipschitz one functions
(3) $\tau_{j}(X_{j})\subset[0,\tau_{max}],$
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed metric space $(X_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance and time preserving timed-Fréchet maps,
(4) $\varphi_{j}\colon X_{j}\to[0,\tau_{max}]\times Z\subset\ell^{\infty},$
where $Z$ is compact, such that
@@
We remark that there are special classes of space-times where the cosmological time functions are easy to control. These include the big-bang space-times and future developed space-times. Sormani–Vega [29] and Sakovich–Sormani [26] defined convergence for such space-times based on pointed GH convergence of Gromov [18] and GH Convergence for Metric Pairs of Che, Galaz-García, Guijarro, and Membrillo Solis [7]. Such notions already have compactness theorems proven by Gromov [18] and by Che and Ahumada Gómez [1] respectively. The intrinsic timed-Hausdorff distance applies to a far wider class of space-times including sub-space-times that exhaust asymptotically flat space-times [26].
It should be noted that some authors have used the Sormani–Vega null distance of [30] to convert space-times into metric spaces without keeping track of the time functions. See the work of Allen, Burtscher, García-Heveling, Kunzinger, and Steinbauer in [2, 3, 6, 20]. Furthermore, Allen, Burtscher, Kunzinger, and Steinbauer [2, 3, 20] have applied Gromov’s original compactness theorem to obtain a GH limit space which is only a metric space with no causal structure unlike our timed metric spaces.
@@
Other notions of weak convergence of space-times do not involve the conversion of the space-times into metric spaces or timed-metric-spaces. Minguzzi–Suhr [22] have a compactness result in the setting of “Lorentzian metric spaces” (which are not metric spaces) building upon work of Noldus [24, 25]. See also the work of Mondino–Sämann [23] in the setting of “Lorentzian pre-length spaces”, as introduced in [19], where they define a notion of Lorentzian Gromov–Hausdorff convergence and obtain a precompactness result using causal diamonds and Lorentzian distances. It is not yet clear how their notions of convergence relate to the Sakovich–Sormani notions defined using timed-metric-spaces in [26]. This is worth exploring further.
@@
We note that, in upcoming work, Sakovich–Sormani will be introducing timed intrinsic flat convergence of space-times whose limit spaces are rectifiable timed-metric-spaces with biLipchitz charts [27]. Our compactness theorem proven in this paper will be applied in that one in an essential way.
The structure of the paper is as follows. In Section 2 we review preliminary notions and results about the Gromov–Hausdorff distance, the classical Gromov’s compactness theorem, and the intrinsic timed-Hausdorff distance. Our compactness theorem is proven in Sections 3.13.3. In Section 3.1 we begin the proof by imitating Gromov’s proof of his compactness theorem in [16]. As we added significant details to his original proof, we developed the notion of “addresses” for points in Section 3.2. With these “addresses” we complete a stronger statement of the compactness theorem which controls the convergence of time in a uniform way in Section 3.3.
Our notion of “addresses” provides a new way of stating Gromov’s Compactness Theorem for sequences of metric spaces, that we state and prove in Section 4.1. We also restate an Arzelà–Ascoli Theorem for sequences of real valued uniformly bounded Lipschitz functions on a converging sequence of metric spaces using “addresses” in Section 4.2. We suggest a more general Arzelà–Ascoli Conjecture that could be proven using “addresses” in Section 4.3.
@@
Acknowledgements: We would like to thank Anna Sakovich (Uppsala) for online conversations in July 2025 that lead us to pursuing this project. We all gratefully acknowledge support from the Simons Center for Geometry and Physics and Stony Brook University where we had the opportunity to meet together in person in September 2025 as part of the Program Geometry and Convergence in Mathematical General Relativity. R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. C. Sormani was funded in part by a PSC-CUNY grant. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ 2.1. Review of the Hausdorff Distance
The Hausdorff distance between subsets, $U,W\subset Z$, lying in a common metric space, $(Z,d_{Z})$, is defined by
(6) $d_{H}^{Z}\Big(U,W\Big)=\inf\left\{r>0:\begin{array}[]{c}\forall u\in U,\exists w\in W\textrm{ s.t. }d_{Z}(u,w)<r\\ \forall w\in W,\exists u\in U\textrm{ s.t. }d_{Z}(u,w)<r\end{array}\right\}.$
Recall that the Hausdorff distance between sets, $U,W\subset Z$, is extrinsic in the sense that it depends on the distance in the extrinsic space.
2.2. Review of Gromov–Hausdorff Distance
Gromov defined his intrinsic Hausdorff distance between metric spaces, $(X_{j},d_{j})$, which do not lie in a common metric space in [18]. This is now called the Gromov–Hausdorff (GH) distance. The GH distance is intrinsic in the sense that it depends only on the intrinsic geometry of the two given metric spaces and not the extrinsic geometry of some common ambient space.
In order to define this distance intrinsically, Gromov considered all possible common metric spaces, $(Z,d_{Z})$, and considered maps: $\varphi_{j}\colon X_{j}\to Z$ that are distance preserving
@@ 2.2. Review of Gromov–Hausdorff Distance
Thanks to the infimum taken over all possible distance preserving maps, the Gromov–Hausdorff distance depends only on the intrinsic geometry of the original pair of metric spaces not on any particular external metric space, $(Z,d_{Z})$. Yet at the same time the GH distance between metric spaces has many of the same properties as the Hausdorff distance. In particular, there are correspondences between $X_{1}$ and $X_{2}$ with controlled distance distortion that are often used to define the GH distance, but we will not use correspondences in this paper.
@@ 2.3. Review of Gromov’s Compactness Theorem
Recall Gromov’s Compactness Theorem proven in [18] and [16]:
@@
If $(X_{j},d_{j})$ is a sequence of compact metric spaces that are equibounded,
@@
then a subsequence converges in the Gromov–Hausdorff sense to a compact metric space $(X_{\infty},d_{\infty})$:
@@
In fact, there exist a compact metric space, $(Z,d_{Z})$, and distance preserving maps,
(12) $\varphi_{j}\colon X_{j}\to Z,$
@@ 2.4. Review of Fréchet Maps
Here we review the notion of a Fréchet map [12].  [^2] In v1 of [26] this was accidentally attributed to Kuratowski. The actual definition of a Kuratowski map can be consulted here [21].
First recall that a separable metric space, $(X,d)$, is a metric space containing a countable dense collection of points.
@@
Given a separable metric space, $(X,d)$, for any countable dense collection of points $\mathcal{N}=\{x_{1},x_{2},$...\} of $X$, one can define a Fréchet map
(16) $\kappa_{X}=\kappa_{X,\mathcal{N}}\colon(X,d)\to(\ell^{\infty},d_{\ell^{\infty}})$
@@
(17) $\kappa_{X}(x)=(d(x_{1},x),d(x_{2},x),\ldots)\subset\ell^{\infty}.$
Now we state Fréchet’s embedding theorem [13]:
@@
It is well known that one can define a version of the Gromov–Hausdorff distance by taking the infimum of the Hausdorff distance between images of Fréchet maps into $\ell^{\infty}$.
@@
The Fréchet-GH distance between two metric spaces is defined as
@@
that is, the infimum is taken over all selections of countably dense points and reorderings of these selected points in $X$ and in $Y$.
By Fréchet’s embedding theorem, for any pair of compact metric-spaces one immediately has
@@
In fact, this notion is biLipschitz equivalent to the Gromov–Hausdorff distance:
(22) $d_{GH}((X,d_{X}),(Y,d_{Y}))\leq d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))\leq 2d_{GH}((X,d_{X}),(Y,d_{Y})).$
This close relationship with the Gromov–Hausdorff distance is discussed and reproven in the background of [26] and directly inspired their definition of the intrinsic timed-Hausdorff distance using timed-Fréchet maps.
@@ 2.5. Review of Timed-Metric-Spaces
The notion of a timed-metric-space was introduced by Sakovich–Sormani in [26] because they had constructed a canonical way to convert certain classes of smooth Lorentzian space-times into timed-metric-spaces keeping track of the causal structure of the original Lorentzian manifold. For our purposes, we need only the abstract notion of a timed-metric-space:
@@
[…]
(23) $\forall q\in X\quad p\in J_{X}^{+}(q)\Longleftrightarrow\tau(p)-\tau(q)=d(p,q).$
Given a map $F\colon(X,d_{X},\tau_{X})\to(Y,d_{Y},\tau_{Y})$ we say that it is distance preserving iff
@@
We say that $F$ preserves the causal structure iff
@@
In this paper we focus on the geometry of timed-metric-spaces and the intrinsic timed-Hausdorff distance. Our work applies to study Lorentzian manifolds and their intrinsic timed-Hausdorff limits, but we will not be exploring this within the paper. Hence, we will not refer further to smooth manifolds.